OpenStudy (anonymous):

Solve for x. 11x + 4 < 37

OpenStudy (anonymous):

x<37/15 x<3 x<41/11 x<363

OpenStudy (anonymous):

think of < as an = sign, because < = and > are just comparisons.

OpenStudy (anonymous):

Your first step would be to remove the +4 on the right hand side. to do this you minus 4 from both sides. the equation should now look like this: 11x + 4 - 4 = 37 - 4

OpenStudy (anonymous):

11x + 4 < 37 Get x by itself. 11x < 33 Divide both sides by 11. x = 3

OpenStudy (anonymous):

of course 4 - 4 = 0 so not the +4 on the right hand side is cancelled out. 11x = 37-4

OpenStudy (anonymous):

I mean x < 3

OpenStudy (muzzack):

11x + 4 < 37 11x<33 x<3

OpenStudy (anonymous):

11x=33 33/11 = 3 x<3

OpenStudy (anonymous):

I think a lot of people have trouble with inequalities because they don't understand that < is just = but it means less than. = means equal, which is a comparison. I try to beat that into heads a lot. :D

OpenStudy (anonymous):

The < in the equation is suppose to have a line under it. Is that the same as equal to?

OpenStudy (anonymous):

Muzzack do you get it

OpenStudy (anonymous):

No, it means less than or equal to. It's the same though.

OpenStudy (anonymous):

It's just a sign.

OpenStudy (anonymous):

x<3

OpenStudy (anonymous):

Ok thanks